OpenStudy (anonymous):

The desert temperature oscillates daily between 4 degrees celsius at 5am and 32 degrees celsius at 5pm. Write a formula for the temperature as a function of time "t" measured in hours from 5am. I know at t = 0, it is 5am. Therefore, at t = 12, it is 5pm. So, (0,4) is a minimum, and (5,32) is a maximum. Drawing just those points out, I can see that it is going to end up as a negative cosine function. I figured that by taking the min and the max y-values and dividing by 2, I get the axis (or the vertical shift), which is 18. Now I am stuck. I have (so far), f(x) = -cos(x) + 18

OpenStudy (anonymous):

no that is not right

OpenStudy (anonymous):

32 - 4 = 24 so you need an amplitude of 12

OpenStudy (anonymous):

if you use \[-12\cos(x)\] then it will fluctuate between -12 and 12

OpenStudy (anonymous):

Wait so a vertical shift of 4 units up is its translation right?

OpenStudy (anonymous):

am i an idiot or what?

OpenStudy (anonymous):

32 - 4 = 28 so you need an amplitude of 14

OpenStudy (anonymous):

is that for saying i am an idiot? thanks ishaan!

OpenStudy (anonymous):

Wait I did not call you an idiot..

OpenStudy (anonymous):

then \[-14\cos(x)\] will fluctuate between -14 and 14 to add 18 like you had

OpenStudy (anonymous):

no i called myself one and ishaan gave me a medal for it!

OpenStudy (anonymous):

so now we should have \[-14\cos(x)+18\] that takes care of the range. now you have to take care of the period

OpenStudy (anonymous):

as it stands if \[x=0\] you get 4 and if \[x=\pi\] you get 32

OpenStudy (anonymous):

Ok so to find the amplitude, it is max - min / 2, to find the axis, it is max + min / 2

OpenStudy (anonymous):

yeah i guess. i wasn't using a formula but it looks like you got it

OpenStudy (anonymous):

OHHH Now I remember.

OpenStudy (anonymous):

Ok so how would you know what the period is, given this information?

OpenStudy (anonymous):

Here, I am confused because you have a min, and you have the ending maximum

OpenStudy (anonymous):

yeah that works because you used \[-\cos(x)\]

OpenStudy (anonymous):

Ok so with the given information, you CANNOT determine a definite period?

OpenStudy (anonymous):

we are going from 4 to 32 right?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

looks like it is 24 hours right?

OpenStudy (anonymous):

5 am for min, 5 pm for max, and then back again

OpenStudy (anonymous):

That is what I was thinking too, but it cannot be.

OpenStudy (anonymous):

why not? we just have to use a 24 hour clock

OpenStudy (anonymous):

so put the period at 24 hours and period is \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

so \[\frac{2\pi}{b}=24\] \[b=\frac{\pi}{12}\]

OpenStudy (anonymous):

and finally we have to adjust to make 5 am 0 so subtract 5

OpenStudy (anonymous):

Ok wait so what is the formula for determining the period?

OpenStudy (anonymous):

period of \[\cos(bx)\] or \[\sin(bx)\] is \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

How do you know that this is the formula? How can you tell that it is right?

OpenStudy (anonymous):

well i know it because i know it, but we can find it by doing an example if you like

OpenStudy (anonymous):

Yes please.

OpenStudy (anonymous):

ok suppose you have \[\sin(6x)\] and you want the period

OpenStudy (anonymous):

Ok can we use the equation that I was taught: Period = 2pi / k

OpenStudy (anonymous):

If you don't mind.

OpenStudy (anonymous):

when \[x=0\] then \[6x=0\] and so we start there. when \[6x=2\pi\] we have completed one full period. solve for x get \[x=\frac{2\pi}{6}\]

OpenStudy (anonymous):

i.e. as x goes from 0 to \[\frac{2\pi}{6}\] 6x goes from 0 to 2pi

OpenStudy (anonymous):

Ok I understand, so you are saying that 6x = 2pi, this is when we have completed a period. But why did you not say sin(6x) = 2pi

OpenStudy (anonymous):

it is an easy enough formula to know and derive if you need it.

OpenStudy (anonymous):

2pi is the input. sin(2pi)=0 is the output

OpenStudy (anonymous):

x= 2pi/6 6x=2pi sin(2pi)=0

OpenStudy (anonymous):

back to the problem at hand since the period is 24 hours we know the period is 24 and therefore \[b=\frac{\pi}{12}\]

OpenStudy (anonymous):

so now we have \[-14\cos(\frac{\pi}{12}x)+18\]

OpenStudy (anonymous):

and finally we have to adjust so that 5 am becomes 0. that is easy. subtract 5 from the input

OpenStudy (anonymous):

\[-14\cos(\frac{\pi}{12}(x-5))+18\]

OpenStudy (anonymous):

what a nice picture! it works. http://www.wolframalpha.com/input/?i=-14cos%28pi%2F12%28x-5%29%29%2B18

OpenStudy (anonymous):

Wait they said STARTING FROM 5AM, so wouldn't that mean that t = 0 at 5am? We wouldn't need to do a shift horizontally..

OpenStudy (anonymous):

i want to put in 5 and take the -cos(0) to get the lowest value

OpenStudy (anonymous):

so i want to shift right 5 units. make 0 into 5

OpenStudy (anonymous):

so yes, like you said, it is a horizontal shift. shift 5 to the right by subtracting by 5

OpenStudy (anonymous):

check the graph and you see that at 5 you get 4, and at 17 (12 hours later) you get 32

OpenStudy (anonymous):

Yea I do not think that you need horizontal shift at all. This is because: f(12) = -14cos(pi/12(12)) + 18 f(12) = 32

OpenStudy (anonymous):

My calculator is in radians.

OpenStudy (anonymous):

we are working in radians. look here http://www.wolframalpha.com/input/?i=-14cos%28pi%2F12%28x-5%29%29%2B18

OpenStudy (anonymous):

you don't need a horizontal shift if you are shifting first in your head. but 12 is not 0 right? we arranged it so if you plug in 5 you get out 4

OpenStudy (anonymous):

But I tried for both (0,4) and (12,32) and it worked without doing a h.shift

OpenStudy (anonymous):

right i agree. but you shifted 5 am to 0 in your head first.

OpenStudy (anonymous):

if you really want 5 to give you 4 you have to shift

OpenStudy (anonymous):

Observe: T(0) = -14cos((pi/2(0)) + 18 = 4 T(12) = -14cos((pi/2(12)) + 18 = 32

OpenStudy (anonymous):

Ok I understand what you mean

OpenStudy (anonymous):

yeah i see. for sure it works but you have said in your head "make 5 am 0 and start counting from there"

OpenStudy (anonymous):

You are saying that if you start at 1am, and you want to move to 5am?

OpenStudy (anonymous):

no i am saying that if you want to make 5 actually give you a temperature of 4, then you have to shift horizontally 5 units right.

OpenStudy (anonymous):

you have "put in 0, get out 4, put in 12, get out 32" which is fine, but if you shift you get "put in 5, get out 4, put in 17, get out 32"

OpenStudy (anonymous):

i agree it is a red herring. we can alway say "start counting at 5 and make that 0"

OpenStudy (anonymous):

OK so then x = 0 will be 12am, from what you are saying?